Let X be a continuous r.v with density function f(x) =(|x| 10 , for −2 < x < 4 0, elsewhere
. Calculate the expected value of X.
1
Expert's answer
2020-05-13T20:23:14-0400

Let "X" be a continuous r.v with density function


"f(x) = \begin{cases}
 {|x|\over 10}, &\text{for}\ -2<x<4 \\
 0 &\text{elsewhere}
\end{cases}"

Calculate the expected value of "X"


"E(X)=\mu_X=\displaystyle\int_{-\infin}^{\infin}xf(x)dx="

"=\displaystyle\int_{-2}^{0}x({-x\over 10})dx+\displaystyle\int_{0}^{4}x({x\over 10})dx="

"=\bigg[-{x^3\over 30}\bigg]\begin{matrix}
 0 \\
 -2
\end{matrix}+\bigg[{x^3\over 30}\bigg]\begin{matrix}
 4 \\
 0
\end{matrix}={8\over 30}+{64\over 30}=2.4"

"E(X)=2.4"